OpenStudy (anonymous):

cos^2x/sin^2x +cosx secx

hartnn (hartnn):

first term : use \(\Large \dfrac{\cos x}{\sin x} = \cot x\) 2nd term, use \(\Large \dfrac{1}{\cos x } = \sec x\)

OpenStudy (anonymous):

can you walk me through the steps please?

OpenStudy (anonymous):

I think Cosx and 1/cosx cancel, right? so is the answer just cotx^2?

hartnn (hartnn):

good start! when we cancel terms form numerator and denominator, '1' remains! \(\dfrac{\cancel a}{\cancel a} =1 \)

hartnn (hartnn):

so, you would get \(\Large \cot^2x+1\) do you know any pythagorean identity in which cot^2 x is present ?

hartnn (hartnn):

do you know \(\Large \csc^2x -\cot^2 x =1 \) ??

OpenStudy (anonymous):

oh!!! Thank you so much! I understand! It is csc^2x :) right?

hartnn (hartnn):

absolutely correct! :) good

hartnn (hartnn):

you seem to be new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

oh, haha, thank you very much!:)(:

hartnn (hartnn):

welcome ^_^